OpenStudy (anonymous):

On which dates of march,2008 will a sunday come?

OpenStudy (zzr0ck3r):

google calendar.....

OpenStudy (anonymous):

tell me how to solve it..??

ganeshie8 (ganeshie8):

lets figure out this year first, then we can go back to 2008 easily

ganeshie8 (ganeshie8):

yesterday is sunday... 06/30 is sunday

ganeshie8 (ganeshie8):

how many days are in each month :- March : April : May : June :

OpenStudy (anonymous):

Are you asking odd days in those months..??

ganeshie8 (ganeshie8):

im asking to fill total number of days in those months

ganeshie8 (ganeshie8):

March : 31 April : 30 May : 31 June : 30

ganeshie8 (ganeshie8):

add total days

OpenStudy (anonymous):

122

ganeshie8 (ganeshie8):

ok, exclude March days from that count.

ganeshie8 (ganeshie8):

June + May + April = 91

ganeshie8 (ganeshie8):

divide it by 7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

13

ganeshie8 (ganeshie8):

91/7 = 13, remainder 0. So, when we go back 3 months, we will end up in a sunday exactly

ganeshie8 (ganeshie8):

03/31 is a sunday

ganeshie8 (ganeshie8):

subtract 7 from 31... and do it until u reach the start of month

OpenStudy (anonymous):

Ok..thanks

ganeshie8 (ganeshie8):

we're not finished yet, wat we doing is for this year

ganeshie8 (ganeshie8):

after doing this, we need to go back 5 years

ganeshie8 (ganeshie8):

since 03/31 is a sunday, 03/24, 03/17, 03/10, 03/03 also would be sundays in 2013

OpenStudy (anonymous):

Its been long procedure...better to calculate odd days for 2008 and added to the month march date..which can exactly divided by 7

ganeshie8 (ganeshie8):

first let me complete my procedure, id like to learn the easy procedure which u saying next... is that fine ?

OpenStudy (anonymous):

Yeah

ganeshie8 (ganeshie8):

Next, having know 03/31 in 2013 is a sunday, you need to go back exactly 5 years :- 5x365 + 1 = 1826 days to land in 03/31 in 2008.

ganeshie8 (ganeshie8):

1826/7 = something, remainder 6

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

so, you need to go back one more day to land in sunday :- 03/30/2008 is a sunday

ganeshie8 (ganeshie8):

now we keep subtracting 7 from 30 :- 03/23, 03/16, 03/09, 03/02

ganeshie8 (ganeshie8):

Overall sundays in march 2008 are :- 03/30, 03/23, 03/16, 03/09, 03/02

ganeshie8 (ganeshie8):

may i knw the easy method you talking about now ?

OpenStudy (zzr0ck3r):

leap year?